An individual only purchases two goods X and Y . The person has an income of $100. The price of X is $2 per

unit and the price of Y is $4 per unit. The utility function of this individual is U = 3X +5Y . This means the person

has a constant marginal utility of 3 and 5 for goods X and Y , respectively. What bundle of goods X and Y should

the person buy to maximize utility. Why? (Answer: 50 units of X and none of Y giving a utility of U = 150)

Please show all steps, and sketch the indifference curve and the budget line representing the optimal bundle of goods.
Answer #1

MRS = MUx/MUy = 3/5 = .6

Px/Py = 2/4 = .5

As MRS > Px/Py

IC is always steeper than budget line

MUx/Px > MUy/Py

Extra dollar spent on gives higher additional Utility, as compared to the, utility when extra dollar spending on Y

so Only X is consumed at eqm

X* = M/Px = 100/2 = 50

Y* = 0

-40- (0, 30) 25 -20- Budget line 20 X (50, 0)eqm